the product of two consecutive positive integers is 288. Find the integers.
Answer by venugopalramana(3286)

the product of two consecutive positive integers is 288. Find the integers.
LET THE 2 NUMBERS BE N,N+1
N(N+1)=288
N^2+N-288=0
NOT POSSIBLE .
THE PROBLEM SHOULD BE PRODUCT OF 2 CONSECUTIVE 'EVEN'INTEGERS IS 288..THEN
ANY INTEGER IS N....ANY EVEN INTEGER IS 2N...NEXT EVEN INTEGER IS 2N+2
2N(2N+2)=288
4N(N+1)=288
N^2+N-72=0
N^2+9N-8N-72=0
N(N+9)-8(N+9)=0
(N-8)(N+9)=0
N=8
SO THE 2 EVEN INTEGERS ARE 2*8=16 AND 16+2=18